UT 20130918 Technical Observing - AGw2 Calibration

Observer: JHill, DMiller (Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: DCox (Tucson)
AO Support: JGuerra (Tucson)
Telescope Support: JLittle (LBTO)
Instrument Support: ESolheid (LBTO)


Testing DX Adaptive Optics at the beginning. They worked for a couple hours, but the heavy clouds rolled over us about midnight.



DX Adaptive Optics Testing



04:00 Closed for humidity. We called Chris and were able to resolve the pointing problem by restarting PSFR.


05:30 Re-open. Sky is partly cloudy. Seeing is about 1.2 arcsec and variable.

07:30 Seeing has improved to 0.8 arcsec FWHM on the guider (even with mirror ventilation off), but thicker clouds are disrupting things.


08:00 Close for continuing heavy clouds.

-- JohnHill - 18 Sep 2013
Topic revision: r2 - 18 Sep 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback